OpenStudy (anonymous):

What is the length of the transverse axis of the hyperbola defined by the equation below? (x-3)^2 (y+5)^2 ________-________=1 4^2 2^2

OpenStudy (anonymous):

-6 would be the legnth?

OpenStudy (anonymous):

so the legnth of the tranverse axis would be 8? could you explain that a little bit more

OpenStudy (dan815):

basically u look at the value the x part is divided by

OpenStudy (dan815):

Given the equation of the hyperbola: [ (y - 36) / (36) ] - [ ( x + 25 ) / (144) ] = 1 Transverse axis is vertical where a = 6, b = 12 Length of the conjugate axis: 2b = 2(12) = 24 Length of the transverse axis: 2a = 2(6) = 12 -------> ANSWER

OpenStudy (dan815):

like that

OpenStudy (anonymous):

got it!

OpenStudy (dan815):

so ur answer is 4

OpenStudy (dan815):

wait i dont know!!, what is the transverse axis

OpenStudy (dan815):



OpenStudy (anonymous):

wait so what you showed me is how to get the legnth ? not the transverse axil? :/